### v1

# The Feldman-Hájek Dichotomy for Countable Gaussian Mixtures and their Asymptotic Separability in High Dimensions

Abstract
This paper establishes the theoretical foundations for the asymptotic separability of Gaussian Mixture Models (GMMs) in high dimensions by extending the classical Feldman-Hájek theorem. We first prove that a countable mixture of Gaussian measures is a well-defined probability measure. Our primary result, the Gaussian Mixture Dichotomy Theorem, demonstrates that the mutual singularity of individual Gaussian components is a sufficient condition for the mutual singularity of the resulting mixtures. We provide a rigorous proof and further discuss the “Mixed Case,” where the presence of even a single equivalent pair of components leads to partial absolute continuity via the Lebesgue decomposition, thereby defining the theoretical limits of perfect classification in infinite-dimensional spaces.
Infinite-Dimensional Measures,
keywords:
[class=MSC]
keywords:
[…]
1 Introduction
The study of probability measures in high-dimensional spaces is often dominated by the phenomenon of measure concentration [2, 3, 4, 5]. As $d\to\infty$ ($d$ being the dimension), the geometry of these measures changes radically, leading to the behavior described for Gaussian measures in 1958 by the Feldman-Hájek dichotomy [4, 5]. This result belongs to a broader class of dichotomy theorems for infinite-dimensional measures, the most notable being Kakutani’s theorem on infinite product measures [7]. The Feldman-Hajék theorem states that two Gaussian measures in an infinite-dimensional Hilbert space are either equivalent or mutually singular. While the behavior of individual Gaussian measures is well-documented, modern statistical learning frequently employs Gaussian Mixture Models (GMMs) to represent complex distributions, and thus it is a relevant questions under which conditions two different GMMs can be perfectly separable in infinite dimensions. Some general discussion about sums of random variables and mixtures can be found in the book by Vakhania et al. in [12], but not a specific result for GMMs.
Several authors have studied GMM models in high dimensional settings, but nobody, to the best knowledge of the author, has given conditions for mutual singularity of GMM models in infinite dimensions. Almost all research work found, dealt with the problem of classifying or learning GMM models, and were not interested in finding conditions for mutually singularity.
It is maybe of interest to the reader to highlight some of the works related to learning GMMs. It should be emphasized that these contributions are, in a certain sense, connected to and conceptually aligned with the research presented in this work. In fact, if two GMMs are mutually singular, it also means that they are perfectly separable, or in other words, in the assumption that one finds the right classifier, one can classify them with zero missclassification error. A first interesting work in this direction, is the one by Mignacco et al. [9]. In their work, they studied the role of regularisation in a classification problem of high-dimensional two-mixture of noisy gaussians. Refinetti et al. studied the problem of classifying high-dimensional Gaussian mixtures, with a focus on neural networks [10] and showed the superiority of neural networks over kernel methods. Azizyan et al. in [1] also studied the problem of learning Gaussian mixtures with equal mixing weights in a high-dimensional settings under small-mean separation. Other works goes along the same line, in studying how to learn gaussian mixtures. The objective of this study is not to address that point directly, but rather to investigate how the Feldman–Hajék theorem can be generalized and extended to the setting of Gaussian mixture measures. As already mentioned, this work is important for the problem of learning mixtures, as it deals with the limit where error-free (perfect) classification becomes theoretically possible.
The primary contribution of this work is the extension of the Feldman-Hájek theorem to countable Gaussian Mixture Models. Specifically, we establish the following. (i) We provide a rigorous proof that a countable mixture of Gaussian measures is a well-defined probability measure on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$. (ii) We prove that the mutual singularity of all constituent Gaussian components is a sufficient condition for the mutual singularity of the resulting mixtures. (iii) We discuss shortly the “mixed case,” demonstrating how the existence of equivalent components breaks the dichotomy and leads to partial absolute continuity. By extending these results to GMMs, we provide a theoretical foundation for the study of asymptotic separability in complex, multi-modal distributions.
The rest of this paper is organised as follows. In Section 2, we establish the formal measure-theoretic framework for countable mixtures and prove that they are valid probability measures via Lemmas 2.1 and 2.2. Section 3 presents our main result, the The Gaussian Mixture Dichotomy Theorem (Theorem 3.1), and its proof. Finally, Section 4 discusses the implications of these results.
2 Formalism and Gaussian Mixtures
Let $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ be a measurable space, where $\mathcal{B}(\mathbb{R}^{d})$ is the Borel $\sigma$-algebra. A countable mixture is a probability measure $\mu$ defined for every $A\subseteq\mathcal{B}(\mathbb{R}^{d})$ as the following.
$\mu(A)=\sum_{k=1}^{\infty}\pi_{k}\gamma_{k}(A)$ (1)
where $\gamma_{k}$ are probability Gaussian measures on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ characterised by the mean vector $m_{k}\in\mathbb{R}^{d}$ and by their covariance $\Sigma_{k}$. $\pi_{k}$ are mixing weights such that $\pi_{k}>0$ and $\sum_{k=1}^{\infty}\pi_{k}=1$. In Lemma 2.2 we prove that $\mu$ itself is a probability measure. Lemma 2.1 is a stepping stone to prove lemma 2.2.
Lemma 2.1.
Given a Gaussian measure $\gamma_{i}$, for any For any sequence of disjoint sets $\{A_{n}\}_{n=1}^{\infty}$ with $A_{n}\subset\mathcal{B}(\mathbb{R}^{d})$ the sum $\sum_{k=1}^{\infty}\gamma_{i}(A_{k})$ exists and is finite.
Proof.
By definition $A_{n}$ are disjoint sets, and since $\gamma_{i}$ is a measure then
$\sum_{n=1}^{\infty}\gamma_{i}(A_{n})=\gamma_{i}\left(\bigcup_{n=1}^{\infty}A_{n}\right)$ (2)
But also by definition $\cup_{n=1}^{\infty}A_{n}\subseteq\mathbb{R}^{d}$ and therefore
$\gamma_{i}\left(\bigcup_{n=1}^{\infty}A_{n}\right)\leq\gamma_{i}(\mathbb{R}^{d})=1$ (3)
Additionally, since each $\gamma_{i}(A_{n})\geq 0$ then the partial sums $s_{m}=\sum_{n=1}^{m}\gamma_{i}(A_{n})$ are monotonically increasing and bounded above by 1, hence it converges. ∎
Lemma 2.2 ($\mu$ is a measure).
Let us define $\mu$ by
$\mu(A)=\sum_{k=1}^{\infty}\pi_{k}\gamma_{k}(A)$ (4)
where $\gamma_{k}$ are probability Gaussian measures on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ characterised by the mean vector $m_{k}\in\mathbb{R}^{d}$ and by their covariance $\Sigma_{k}$. $\pi_{k}$ are mixing weights such that $\pi_{k}>0$ and $\sum_{k=1}^{\infty}\pi_{k}=1$. $\mu$ is a probability measure.
Proof.
To prove that $\mu$ is a measure we need to prove three results from the definition of measure [11].
  1. (i)
    Non-negativity. $\mu(A)\geq 0$ since all $\pi_{k}>0$ and, since they are measures, $\gamma_{k}(A)\geq 0$.
  2. (ii)
    Empty set. $\mu(\emptyset)=\sum_{k=1}^{\infty}\pi_{k}\gamma(\emptyset)=\sum_{k=1}^{\infty}\pi_{k}\cdot 0=0$.
  3. (iii)
    $\sigma$-additivity. For any sequence of disjoint sets $\{A_{n}\}_{n=1}^{\infty}$, we can write
    $\mu\left(\bigcup_{n=1}^{\infty}A_{n}\right)=\sum_{i=1}^{\infty}\pi_{i}\gamma_{i}\left(\bigcup_{n=1}^{\infty}A_{n}\right)=\sum_{i=1}^{\infty}\pi_{i}\sum_{n=1}^{\infty}\gamma_{i}(A_{n})$ (5)
    By Lemma 2.1, since the second sum is finite (the partial sums converge), we can exchange the two sums, giving
    $\mu\left(\bigcup_{n=1}^{\infty}A_{n}\right)=\sum_{n=1}^{\infty}\sum_{i=1}^{\infty}\pi_{i}\gamma_{i}(A_{n})=\sum_{n=1}^{\infty}\mu(A_{n})$ (6)
    concluding the proof.
Now let us consider two GMMs $\mu$ and $\nu$ defined over two disjoint sets of indices $I_{1}$ and $I_{2}$ similarly to Equation 1. Namely we will define
$\mu$ $=\sum_{i\in I_{1}}\pi_{i}\gamma_{i}$ (7)
$\nu$ $=\sum_{i\in I_{2}}\pi_{i}\gamma_{i}$ (8)
Now we will prove the following lemma.
Lemma 2.3.
If $\gamma_{i}\perp\gamma_{j}$ (the two measures are mutually singular) $\forall i\in I_{1}$ and $\forall j\in I_{2}$ then $\mu\perp\nu$.
Proof.
To prove this lemma, we need to show that $\exists A$ such that $\mu(A)=1$ and $\nu(A)=0$. Let us construct such a set. Recall that, by assumption, $\gamma_{i}\perp\gamma_{j}$ (the two measures are mutually singular) $\forall i\in I_{1}$ and $\forall j\in I_{2}$, thus there exists a set $A_{ij}$ such that $\gamma_{i}(A_{ij})=1$ and $\gamma_{j}(A_{ij})=0$. Consider the set
$A=\bigcup_{i\in I_{1}}\left(\bigcap_{j\in I_{2}}A_{ij}\right)$ (10)
We want to show that $\mu(A)=1$ and $\nu(A)=0$, thus proving the lemma. To simplify our notation, let us define
$S_{i}=\bigcap_{j\in I_{2}}A_{ij}$ (11)
First of all, let us show that $\gamma_{j}(S_{i})=0$ for $j\in I_{2}$ and $i\in I_{1}$. In fact
$S_{i}=\bigcap_{j\in I_{2}}A_{ij}\subseteq\bigcup_{j\in I_{2}}A_{ij}$ (12)
therefore
$\gamma_{j}(S_{i})\leq\gamma_{j}\left(\bigcup_{j\in I_{2}}A_{ij}\right)=\sum_{j\in I_{2}}\gamma_{j}(A_{ij})=0$ (13)
Now we can prove that $\gamma_{i}(S_{i})=1$ for $j\in I_{2}$ and $i\in I_{1}$. To prove this it is enough to prove that $\gamma_{i}(S_{i}^{c})=0$ since $\gamma_{i}$ is a probability measure. Thanks to the De Morgan’s law
$S_{i}^{c}=\left(\bigcap_{j\in I_{2}}A_{ij}\right)^{c}=\bigcup_{j\in I_{2}}A_{ij}^{c}$ (14)
it follows
$\gamma_{i}(S_{i}^{c})\leq\sum_{j\in I_{2}}\gamma_{j}(A_{ij}^{c})=0$ (15)
since the $A_{ij}^{c}$ are not necessarly mutually disjoint so we use the $\sigma$-subadditivity property of $\gamma_{j}$. So we established that $\gamma_{i}(S_{i})=1$ for $i\in I_{1}$ and $\gamma_{j}(S_{i})=0$ for $j\in I_{2}$ and $i\in I_{1}$. Now let us turn our attention to $A$ defined in Equation (10).
$\nu(A)=\nu\left(\bigcup_{i\in I_{1}}\left(\bigcap_{j\in I_{2}}A_{ij}\right)\right)\leq\sum_{i\in I_{1}}\nu\left(\bigcap_{j\in I_{2}}A_{ij}\right)=\sum_{i\in I_{1}}\sum_{j\in I_{2}}\pi_{j}\gamma_{j}(S_{i})=0$ (16)
since we have proved that $\gamma_{j}(S_{i})=0$. Thus we have proven that $\nu(A)=0$. Now let us prove that $\mu(A)=1$. We have
$A=\bigcup_{i\in I_{1}}S_{i}\Rightarrow A^{c}=\bigcap_{i\in I_{1}}S_{i}^{c}$ (17)
Now
$\bigcup_{i\in I_{1}}S_{i}\subseteq\mathbb{R}^{d}\Rightarrow\mu\left(\bigcup_{i\in I_{1}}S_{i}\right)\leq\mu(\mathbb{R}^{d})=1$ (18)
But at the same time
$S_{1}\subseteq\bigcup_{i\in I_{1}}S_{i}\Rightarrow 1=\mu(S_{1})\leq\mu\left(\bigcup_{i\in I_{1}}S_{i}\right)$ (19)
Or in other words we have proven that $1\leq\mu(A)\leq 1$ and therefore $\mu(A)=1$. Thus we have proven that $\mu\perp\nu$. ∎
Now we can give our main results of this paper: the extension of the Feldman-Hajek theoreom for GMMs.
3 The Gaussian Mixture Dichotomy Theorem
Based on the construction of the measure $\mu$ as a countable mixture of Gaussian measures the singularity result established in Lemma 2.3, we state the main result.
Theorem 3.1 (The Gaussian Mixture Dichotomy Theorem).
Let $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ be a measurable space. Let $\mu$ and $\nu$ be two Gaussian Mixture Models (GMMs) defined over disjoint index sets $I_{1}$ and $I_{2}$ as
$\mu=\sum_{i\in I_{1}}\pi_{i}\gamma_{i},\quad\nu=\sum_{j\in I_{2}}\rho_{j}\gamma_{j}$ (20)
where $\gamma_{k}=\mathcal{N}(m_{k},\Sigma_{k})$ are Gaussian measures with mean vectors $m_{k}$ and covariance $\Sigma_{k}$. The measures $\mu$ and $\nu$ are mutually singular ($\mu\perp\nu$) if for every pair $(i,j)\in I_{1}\times I_{2}$, the measures $\gamma_{i}$ and $\gamma_{j}$ satisfy the Feldman-Hájek singularity conditions.
Specifically, $\mu\perp\nu$ if for all $i\in I_{1}$ and $j\in I_{2}$, at least one of the following conditions is met:
  1. 1.
    If
    $T=\Sigma_{i}^{-1/2}\Sigma_{j}\Sigma_{i}^{-1/2}-I$ (21)
    is not a Hilbert-Schmidt operator, i.e.,
    $\sum_{n=1}^{\infty}\lambda_{n}^{2}=\infty$ (22)
    where $\lambda_{n}$ are the eigenvalues of $T$.
  2. 2.
    The mean difference $(m_{i}-m_{j})$ does not belong to the common Cameron-Martin Space $H=\Sigma_{i}^{1/2}(\mathbb{R}^{d})=\Sigma_{j}^{1/2}(\mathbb{R}^{d})$.
Proof.
By Lemma 2.3, if $\gamma_{i}\perp\gamma_{j}$ for all $i\in I_{1}$ and $j\in I_{2}$, then $\mu\perp\nu$. Under the standard Feldman-Hájek theorem, two Gaussian measures are mutually singular if and only if the aforementioned operator or mean conditions are violated. Since these conditions imply $\gamma_{i}\perp\gamma_{j}$, the singularity of the mixtures follows from the existence of the set $A=\bigcup_{i\in I_{1}}(\bigcap_{j\in I_{2}}A_{ij})$ such that $\mu(A)=1$ and $\nu(A)=0$. ∎
4 Results and Discussion
The extension of the Feldman-Hájek theorem to countable Gaussian mixtures establishes a rigorous framework for understanding the separability of complex distributions in high dimensions. A critical nuance of this result is its application across both finite and infinite-dimensional Hilbert spaces.
In finite dimensions ($\mathbb{R}^{d}$), two non-degenerate Gaussian measures are almost always equivalent because their supports coincide with the entire space. However, singularity ($\gamma_{i}\perp\gamma_{j}$) can occur in finite dimensions if the measures are degenerate, for instance, if they are supported on disjoint lower-dimensional affine subspaces.
The The Gaussian Mixture Dichotomy Theorem becomes interesting as $d\to\infty$. In infinite-dimensional spaces, the dichotomy takes hold: two Gaussian measures are either equivalent or mutually singular. This rigorous disjointness of support in infinite dimensions is a characteristic property of Gaussian measures [4] and is in general valid for product-type measures [7] and, as we have demonstrated, scales to Gaussian Mixture Models (GMMs). If the components of the mixtures are singular, a condition almost always satisfied in the infinite-dimensional limit due to shifts in mean or scaling of covariance, the global mixtures $\mu$ and $\nu$ are themselves perfectly separable.
4.1 Implications for Statistical Learning
The proof that GMMs can be (under certain assumptions) perfectly separable in the limit of infinite dimensions implies the existence of a perfect classifier in the limit. This provides a theoretical ceiling for Bayesian inference and classification tasks involving Gaussian processes and their mixtures, confirming that the complexity of the mixture (even if countable) does not degrade the inherent separability provided by the Feldman-Hájek conditions.
4.2 The Mixed Case
An important aspect in applying the The Gaussian Mixture Dichotomy Theorem arises when the index sets $I_{1}$ and $I_{2}$ each include Gaussian pairs that are not all of the same type, but instead comprise a combination of pairs that are equivalent and pairs that are mutually singular. While the standard theorem for pure Gaussian measures permits only a binary state (equivalence or singularity), the mixture model $\mu$ and $\nu$ allows for a ”mixed” or intermediate state.
If there exists at least one pair $(i,j)\in I_{1}\times I_{2}$ such that $\gamma_{i}\sim\gamma_{j}$, the perfect separability established in Lemma 2.3 is lost. In this case, the measures $\mu$ and $\nu$ are no longer mutually singular. Instead, they exhibit partial absolute continuity, which can be formally described via the Lebesgue Decomposition [6, 11]:
$\nu=\nu_{ac}+\nu_{s}$ (23)
where $\nu_{ac}$ is the portion of the mixture formed by components that share support with $\mu$, and $\nu_{s}$ represents the components that remain singular to all $\gamma_{i}$ for $i\in I_{1}$.
In practical terms, this means that the strictly “all-or-nothing” nature of separability in GMMs fundamentally depends on the pair-wise singularity condition being satisfied by their individual components. If even a single pair of components in the two mixtures is equivalent, the intersection of their supports creates a region of support overlap for both measures, precluding the existence of an error-free decision boundary. This underscores the importance of the condition $\gamma_{i}\perp\gamma_{j}\,\forall i\in I_{1},j\in I_{2}$ in high-dimensional settings.
5 Practical Applications
The Gaussian Mixture Dichotomy Theorem (Theorem 3.1) provides a rigorous mathematical foundation for the phenomenon of “perfect classification” in high-dimensional settings. While the original theorem is restricted to single Gaussian processes [4, 5], real-world datasets, such as high-resolution images, genomic sequences, spectral data, or financial time series, are often characterised by multi-modality and are effectively modeled by Gaussian Mixture Models (GMMs).
In computer science and signal processing, a primary objective is the error-free separation of classes. Theorem 3.1 establishes that if these classes are modelled by GMMs whose components satisfy the Feldman-Hájek singularity conditions, a decision boundary exists that achieves a zero error rate in the infinite-dimensional limit. This provides a theoretical ceiling for the performance of statistical learning algorithms, confirming that the multi-modal complexity of the mixture does not prevent perfect separability as long as the underlying components are mutually singular. The result shows how pure measure theory revelas itself in applied machine learning by demonstrating that the “all-or-nothing” dichotomy remains robust even for countable mixtures.
Since GMMs are universal approximators for continuous probability densities, the implications of Lemma 2.3 extend to nearly any dataset that can be modelled as a mixture of localised concentrations of measure. In high-dimensional anomaly detection, the theorem predicts that an anomaly, when modelled as a singular mixture component, will not disappear into the baseline data distribution as the dimensionality $d\to\infty$. Instead, the concentration of measure on thin shells ensures that these mixtures remain perfectly distinguishable, as their respective supports are disjoint in the infinite dimensional limit.
The author successfully applied this theorem and verified its validity in spectroscopy data (fluorescence spectroscopy of olive oil) and found that in real data, it is possible to reach an extremely high accuracy in classification even from spectral regions that do not contain any chemical information relevant to the classification [8] given a sufficient number of input spectral intensities.
For real-world data where complete singularity is not present, the analysis of the “Mixed Case” allows researchers to quantify the inherent “error floor” of a system. If data analysis reveals that certain components of a GMM are equivalent rather than singular, the Lebesgue decomposition identifies the specific portion of the data that is mathematically inseparable. This allows to distinguish between classification errors caused by sub-optimal model selection and errors that are fundamentally mandatory due to the presence of equivalent probability measures (assuming the number of dimensions is high enough).
{acks}
[Acknowledgments] The author would like to thank Prof. C. Schwab, that pointed out the Feldman-Hajék theorem to me, during a workshop on infinite dimensional statistics.

### v2

@@
-The Feldman-Hájek Dichotomy for Countable Gaussian Mixtures and their Asymptotic Separability in High Dimensions
-Abstract
-This paper establishes the theoretical foundations for the asymptotic separability of Gaussian Mixture Models (GMMs) in high dimensions by extending the classical Feldman-Hájek theorem.
-We first prove that a countable mixture of Gaussian measures is a well-defined probability measure.
-Our primary result, the Gaussian Mixture Dichotomy Theorem, demonstrates that the mutual singularity of individual Gaussian components is a sufficient condition for the mutual singularity of the resulting mixtures.
-We provide a rigorous proof and further discuss the “Mixed Case,” where the presence of even a single equivalent pair of components leads to partial absolute continuity via the Lebesgue decomposition, thereby defining the theoretical limits of perfect classification in infinite-dimensional spaces.
-Infinite-Dimensional Measures,
-keywords:
-[class=MSC]
-keywords:
+
+[…]
 […]
@@
 The study of probability measures in high-dimensional spaces is often dominated by the phenomenon of measure concentration [2, 3, 4, 5].
-As $d\to\infty$ ($d$ being the dimension), the geometry of these measures changes radically, leading to the behavior described for Gaussian measures in 1958 by the Feldman-Hájek dichotomy [4, 5].
+As one transitions to infinite-dimensional settings, the geometry of these measures changes radically, leading to the behavior described for Gaussian measures in 1958 by the Feldman-Hájek dichotomy [4, 5].
 This result belongs to a broader class of dichotomy theorems for infinite-dimensional measures, the most notable being Kakutani’s theorem on infinite product measures [7].
@@
 Several authors have studied GMM models in high dimensional settings, but nobody, to the best knowledge of the author, has given conditions for mutual singularity of GMM models in infinite dimensions.
-Almost all research work found, dealt with the problem of classifying or learning GMM models, and were not interested in finding conditions for mutually singularity.
-It is maybe of interest to the reader to highlight some of the works related to learning GMMs.
-It should be emphasized that these contributions are, in a certain sense, connected to and conceptually aligned with the research presented in this work.
-In fact, if two GMMs are mutually singular, it also means that they are perfectly separable, or in other words, in the assumption that one finds the right classifier, one can classify them with zero missclassification error.
+Almost all research work found dealt with the problem of classifying or learning GMM models, and were not interested in finding conditions for mutual singularity.
+It may be of interest to the reader to highlight some of the works related to learning GMMs.
+It should be emphasised that these contributions are in a certain sense connected to and conceptually aligned with the research presented in this work.
+In fact, if two GMMs are mutually singular, it also means that they are perfectly separable, or in other words, under the assumption that one finds the right classifier, one can classify them with zero missclassification error.
 A first interesting work in this direction, is the one by Mignacco et al.
@@
 Other works goes along the same line, in studying how to learn gaussian mixtures.
-The objective of this study is not to address that point directly, but rather to investigate how the Feldman–Hajék theorem can be generalized and extended to the setting of Gaussian mixture measures.
+The objective of this study is not to address that point directly, but rather to investigate how the Feldman–Hajék theorem can be generalized and extended to the setting of countable Gaussian mixture measures.
 As already mentioned, this work is important for the problem of learning mixtures, as it deals with the limit where error-free (perfect) classification becomes theoretically possible.
@@
 Specifically, we establish the following.
-(i) We provide a rigorous proof that a countable mixture of Gaussian measures is a well-defined probability measure on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$.
-(ii) We prove that the mutual singularity of all constituent Gaussian components is a sufficient condition for the mutual singularity of the resulting mixtures.
-(iii) We discuss shortly the “mixed case,” demonstrating how the existence of equivalent components breaks the dichotomy and leads to partial absolute continuity.
-By extending these results to GMMs, we provide a theoretical foundation for the study of asymptotic separability in complex, multi-modal distributions.
+(i) We provide a rigorous proof that a countable mixture of Gaussian measures is a well-defined probability measure.
+(ii) We prove that the mutual singularity of all Gaussian components of the mixtures is a sufficient condition for the mutual singularity of the resulting mixtures.
+(iii) We discuss briefly the “mixed case” demonstrating how the existence of equivalent components breaks the dichotomy and leads to partial absolute continuity.
+By extending these results to GMMs, we provide the first steps for the theoretical foundation for the study of asymptotic separability in complex, multi-modal distributions.
 The rest of this paper is organised as follows.
-In Section 2, we establish the formal measure-theoretic framework for countable mixtures and prove that they are valid probability measures via Lemmas 2.1 and 2.2.
-Section 3 presents our main result, the The Gaussian Mixture Dichotomy Theorem (Theorem 3.1), and its proof.
+In Section 2, we establish the formal measure-theoretic framework for countable mixtures and prove that they are valid probability measures via Lemmas 2.1 and 2.3.
+Section 3 presents our main result, the Gaussian Mixture Dichotomy Theorem (Theorem 3.1), and its proof.
 Finally, Section 4 discusses the implications of these results.
 2 Formalism and Gaussian Mixtures
-Let $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ be a measurable space, where $\mathcal{B}(\mathbb{R}^{d})$ is the Borel $\sigma$-algebra.
-A countable mixture is a probability measure $\mu$ defined for every $A\subseteq\mathcal{B}(\mathbb{R}^{d})$ as the following.
+Let $(X,\mathcal{B}(X))$ be a measurable space, where $\mathcal{B}(X)$ is a $\sigma$-algebra.
+A countable mixture is a probability measure $\mu$ defined for every $A\subseteq\mathcal{B}(X)$ as the following.
 $\mu(A)=\sum_{k=1}^{\infty}\pi_{k}\gamma_{k}(A)$ (1)
-where $\gamma_{k}$ are probability Gaussian measures on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ characterised by the mean vector $m_{k}\in\mathbb{R}^{d}$ and by their covariance $\Sigma_{k}$.
-$\pi_{k}$ are mixing weights such that $\pi_{k}>0$ and $\sum_{k=1}^{\infty}\pi_{k}=1$.
+where $\gamma_{k}$ are probability Gaussian measures on $(X,\mathcal{B}(X))$ characterised by the mean vector $m_{k}$ and by their covariance $\Sigma_{k}$.
+$\pi_{k}$ are mixing weights such that $\pi_{k}\geq 0$ and $\sum_{k=1}^{\infty}\pi_{k}=1$.
 In Lemma 2.3 we prove that $\mu$ itself is a probability measure.
@@
 Lemma 2.1.
-Given a Gaussian measure $\gamma_{i}$, for any For any sequence of disjoint sets $\{A_{n}\}_{n=1}^{\infty}$ with $A_{n}\subset\mathcal{B}(\mathbb{R}^{d})$ the sum $\sum_{k=1}^{\infty}\gamma_{i}(A_{k})$ exists and is finite.
-Proof.
+Given a Gaussian measure $\gamma_{i}$, for any For any sequence of disjoint sets $\{A_{n}\}_{n=1}^{\infty}$ with $A_{n}\subset(X,\mathcal{B}(X))$ the sum $\sum_{k=1}^{\infty}\gamma_{i}(A_{k})$ exists and is finite.
+Proof 2.2.
 By definition $A_{n}$ are disjoint sets, and since $\gamma_{i}$ is a measure then
 $\sum_{n=1}^{\infty}\gamma_{i}(A_{n})=\gamma_{i}\left(\bigcup_{n=1}^{\infty}A_{n}\right)$ (2)
-But also by definition $\cup_{n=1}^{\infty}A_{n}\subseteq\mathbb{R}^{d}$ and therefore
-$\gamma_{i}\left(\bigcup_{n=1}^{\infty}A_{n}\right)\leq\gamma_{i}(\mathbb{R}^{d})=1$ (3)
-Additionally, since each $\gamma_{i}(A_{n})\geq 0$ then the partial sums $s_{m}=\sum_{n=1}^{m}\gamma_{i}(A_{n})$ are monotonically increasing and bounded above by 1, hence it converges. ∎
-Lemma 2.2 ($\mu$ is a measure).
+But also by definition $\cup_{n=1}^{\infty}A_{n}\subseteq X$ and therefore
+$\gamma_{i}\left(\bigcup_{n=1}^{\infty}A_{n}\right)\leq\gamma_{i}(X)=1$ (3)
+Additionally, since each $\gamma_{i}(A_{n})\geq 0$ then the partial sums $s_{m}=\sum_{n=1}^{m}\gamma_{i}(A_{n})$ are monotonically increasing and bounded above by 1, hence the partial sums converge.
+This concludes the proof.
+Lemma 2.3 ($\mu$ is a probability measure).
 Let us define $\mu$ by
 $\mu(A)=\sum_{k=1}^{\infty}\pi_{k}\gamma_{k}(A)$ (4)
-where $\gamma_{k}$ are probability Gaussian measures on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ characterised by the mean vector $m_{k}\in\mathbb{R}^{d}$ and by their covariance $\Sigma_{k}$.
-$\pi_{k}$ are mixing weights such that $\pi_{k}>0$ and $\sum_{k=1}^{\infty}\pi_{k}=1$.
+where $\gamma_{k}$ are probability Gaussian measures on $(X,\mathcal{B}(X))$ characterised by the mean vector $m_{k}$ and by their covariance $\Sigma_{k}$.
+$\pi_{k}$ are mixing weights such that $\pi_{k}\geq 0$ and $\sum_{k=1}^{\infty}\pi_{k}=1$.
 $\mu$ is a probability measure.
-Proof.
+Proof 2.4.
 To prove that $\mu$ is a measure we need to prove three results from the definition of measure [11].
@@
     concluding the proof.
-Now let us consider two GMMs $\mu$ and $\nu$ defined over two disjoint sets of indices $I_{1}$ and $I_{2}$ similarly to Equation 1. Namely we will define
+Now let us consider two GMMs $\mu$ and $\nu$ defined over two disjoint sets of indices $I_{1}$ and $I_{2}$ similarly to Equation 1. Now we will prove the following lemma.
+Lemma 2.5.
+Consider two disjoint set of indeces $I_{1}$ and $I_{2}$.
+Given $\gamma_{i}\ i\in I_{1}$ and $\gamma_{j}\ j\in I_{2}$ gaussian measures such that $\gamma_{i}\perp\gamma_{j}$ (the two measures are mutually singular) $\forall i\in I_{1}$ and $\forall j\in I_{2}$.
+Defining
 $\mu$ $=\sum_{i\in I_{1}}\pi_{i}\gamma_{i}$ (7)
 $\nu$ $=\sum_{i\in I_{2}}\pi_{i}\gamma_{i}$ (8)
-Now we will prove the following lemma.
-Lemma 2.3.
-If $\gamma_{i}\perp\gamma_{j}$ (the two measures are mutually singular) $\forall i\in I_{1}$ and $\forall j\in I_{2}$ then $\mu\perp\nu$.
-Proof.
+With $\pi_{i}\geq 0$ and $\sum_{i=1}^{\infty}\pi_{i}=1$. it follows that $\mu\perp\nu$.
+Proof 2.6.
 To prove this lemma, we need to show that $\exists A$ such that $\mu(A)=1$ and $\nu(A)=0$.
 Let us construct such a set.
-Recall that, by assumption, $\gamma_{i}\perp\gamma_{j}$ (the two measures are mutually singular) $\forall i\in I_{1}$ and $\forall j\in I_{2}$, thus there exists a set $A_{ij}$ such that $\gamma_{i}(A_{ij})=1$ and $\gamma_{j}(A_{ij})=0$.
+Recall that, by assumption, $\gamma_{i}\perp\gamma_{j}$ (the two measures are mutually singular) $\forall i\in I_{1}$ and $\forall j\in I_{2}$, thus there exists a set $A_{ij}$ such that $\gamma_{i}(A_{ij})=1$ and $\gamma_{j}(A_{ij})=0$ $\forall i\in I_{1}$ and $\forall j\in I_{2}$.
 Consider the set
@@
 Now
-$\bigcup_{i\in I_{1}}S_{i}\subseteq\mathbb{R}^{d}\Rightarrow\mu\left(\bigcup_{i\in I_{1}}S_{i}\right)\leq\mu(\mathbb{R}^{d})=1$ (18)
+$\bigcup_{i\in I_{1}}S_{i}\subseteq X\Rightarrow\mu\left(\bigcup_{i\in I_{1}}S_{i}\right)\leq\mu(X)=1$ (18)
 But at the same time
 $S_{1}\subseteq\bigcup_{i\in I_{1}}S_{i}\Rightarrow 1=\mu(S_{1})\leq\mu\left(\bigcup_{i\in I_{1}}S_{i}\right)$ (19)
-Or in other words we have proven that $1\leq\mu(A)\leq 1$ and therefore $\mu(A)=1$. Thus we have proven that $\mu\perp\nu$. ∎
+Or in other words we have proven that $1\leq\mu(A)\leq 1$ and therefore $\mu(A)=1$. Thus we have proven that $\mu\perp\nu$.
 Now we can give our main results of this paper: the extension of the Feldman-Hajek theoreom for GMMs.
@@
 Theorem 3.1 (The Gaussian Mixture Dichotomy Theorem).
-Let $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ be a measurable space. Let $\mu$ and $\nu$ be two Gaussian Mixture Models (GMMs) defined over disjoint index sets $I_{1}$ and $I_{2}$ as
+Let $(X,\mathcal{B}(X))$ be a measurable space. Let $\mu$ and $\nu$ be two Gaussian Mixture Models (GMMs) defined over disjoint index sets $I_{1}$ and $I_{2}$ as
 $\mu=\sum_{i\in I_{1}}\pi_{i}\gamma_{i},\quad\nu=\sum_{j\in I_{2}}\rho_{j}\gamma_{j}$ (20)
@@
   2. 2.
-    The mean difference $(m_{i}-m_{j})$ does not belong to the common Cameron-Martin Space $H=\Sigma_{i}^{1/2}(\mathbb{R}^{d})=\Sigma_{j}^{1/2}(\mathbb{R}^{d})$.
-Proof.
+    The mean difference $(m_{i}-m_{j})$ does not belong to the common Cameron-Martin Space $H=\Sigma_{i}^{1/2}(X)=\Sigma_{j}^{1/2}(X)$.
+Proof 3.2.
 By Lemma 2.3, if $\gamma_{i}\perp\gamma_{j}$ for all $i\in I_{1}$ and $j\in I_{2}$, then $\mu\perp\nu$.
@@
 Since these conditions imply $\gamma_{i}\perp\gamma_{j}$, the singularity of the mixtures follows from the existence of the set $A=\bigcup_{i\in I_{1}}(\bigcap_{j\in I_{2}}A_{ij})$ such that $\mu(A)=1$ and $\nu(A)=0$.
-∎
 4 Results and Discussion
@@
 A critical nuance of this result is its application across both finite and infinite-dimensional Hilbert spaces.
-In finite dimensions ($\mathbb{R}^{d}$), two non-degenerate Gaussian measures are almost always equivalent because their supports coincide with the entire space.
-However, singularity ($\gamma_{i}\perp\gamma_{j}$) can occur in finite dimensions if the measures are degenerate, for instance, if they are supported on disjoint lower-dimensional affine subspaces.
-The The Gaussian Mixture Dichotomy Theorem becomes interesting as $d\to\infty$.
+In finite dimensions (e.g.
+$\mathbb{R}^{d}$), two non-degenerate Gaussian measures are almost always equivalent because their supports coincide with the entire space.
+However, singularity ($\gamma_{i}\perp\gamma_{j}$) can occur in finite dimensions if the measures are degenerate, for instance, or if they are supported on disjoint lower-dimensional affine subspaces.
+The Gaussian Mixture Dichotomy Theorem becomes particularly relevant in infinite-dimensional spaces.
 In infinite-dimensional spaces, the dichotomy takes hold: two Gaussian measures are either equivalent or mutually singular.
@@
 4.2 The Mixed Case
-An important aspect in applying the The Gaussian Mixture Dichotomy Theorem arises when the index sets $I_{1}$ and $I_{2}$ each include Gaussian pairs that are not all of the same type, but instead comprise a combination of pairs that are equivalent and pairs that are mutually singular.
+An important aspect in applying the Gaussian Mixture Dichotomy Theorem arises when the index sets $I_{1}$ and $I_{2}$ each include Gaussian pairs that are not all of the same type, but instead comprise a combination of pairs that are equivalent and pairs that are mutually singular.
 While the standard theorem for pure Gaussian measures permits only a binary state (equivalence or singularity), the mixture model $\mu$ and $\nu$ allows for a ”mixed” or intermediate state.
@@
 If even a single pair of components in the two mixtures is equivalent, the intersection of their supports creates a region of support overlap for both measures, precluding the existence of an error-free decision boundary.
-This underscores the importance of the condition $\gamma_{i}\perp\gamma_{j}\,\forall i\in I_{1},j\in I_{2}$ in high-dimensional settings.
+This highlight the importance of the condition $\gamma_{i}\perp\gamma_{j}\,\forall i\in I_{1},j\in I_{2}$ in high-dimensional settings.
 5 Practical Applications
@@
 Since GMMs are universal approximators for continuous probability densities, the implications of Lemma 2.5 extend to nearly any dataset that can be modelled as a mixture of localised concentrations of measure.
-In high-dimensional anomaly detection, the theorem predicts that an anomaly, when modelled as a singular mixture component, will not disappear into the baseline data distribution as the dimensionality $d\to\infty$.
+In high-dimensional anomaly detection, the theorem predicts that an anomaly, when modelled as a singular mixture component, will not disappear into the baseline data distribution in the infinite-dimensional limit.
 Instead, the concentration of measure on thin shells ensures that these mixtures remain perfectly distinguishable, as their respective supports are disjoint in the infinite dimensional limit.
@@
 This allows to distinguish between classification errors caused by sub-optimal model selection and errors that are fundamentally mandatory due to the presence of equivalent probability measures (assuming the number of dimensions is high enough).
-{acks}
-[Acknowledgments] The author would like to thank Prof. C. Schwab, that pointed out the Feldman-Hajék theorem to me, during a workshop on infinite dimensional statistics.